Grade 12AIPMT

THE WORK DONE IN SLOWLY LIFTING A BODY FROM SURFACE TO HEIGHT r IS TWICE THE AMT. OF WORK DONE IN LIFTING THE SAME BODY FROM EARTH SURFACE TO HEIGHT h . FIND THE HEIGHT h.

RAGINI SEIWAL
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
Dear student
 
work done will be stored as potential energy
2[ -GMm/(R+h) + GMm/R ]= -GMm/2R + GMm/R
3GMm/4R = GMm/(R+h)
3h = 4R
h = R/3
 
Regards
Arun (askIITians forum expert)